Suppose that the price of commodity Y is $1 per unit while the price of commodity X is $2 per unit and suppose that an individual’s money income is $16 per time period and is all spent on X and Y. (a) Draw the budget constraint line for this consumer and (b) explain the reason for the shape and the properties of the budget constraint line in part (a).


1
Expert's answer
2021-12-08T19:31:02-0500

The budget constraint usually takes the form

"M=Px+Py"

Where :

M=Income

P=Price

x and y are the products.

In ths case, the budget constraint will be,

"16=2X+Y"

a. To draw the budget constraint line, we determine both X and Y intercepts

"16=2X+Y"

"X=8-0.5Y"

Therefore X intercept is 8

"16=2X+Y"

"Y=16-2X"

and Y intercept is 16

The budget constraint for the consumer is therefore a staright line cutting through Y at 16 and X at 8.

b. The budget constraint equation is a linear equation therefore justifying the linear shape of the budget constraint line. The slope of the graph is is "-0.5" implying that product X is twice as costly as product Y.